# MOCHA: Multi-modal Objects-aware Cross-arcHitecture Alignment

### v3

@@ 4.2 Analyses
Fig. 4 evaluates the role of teacher supervision dimensionality ($d_{t}$) by comparing features: (i) of personal datasets $\mathcal{D}_{f}$ (as an upper bound), (ii) of the distillation dataset $\mathcal{D}_{c}$ unnormalized and (iii) PCA normalized (additional details and validation in Appendix). Overall, performance increases steadily up to $d_{t}=512$, which strikes a good balance between compactness and expressivity. After that, it plateaus or slightly degrades up to the maximum dimension $d=4608$. This indicates that most of the relevant semantics are preserved in a compact subspace, and that further dimensions may introduce redundancy or noise. Normalized features $u_{i}^{\prime}$ consistently outperform their unnormalized counterparts $\hat{u}_{i}^{\prime}$ (especially in the 1-shot setting), obtaining better performance even if compared to personal datasets’ features.
YOLO
@@
Figure 7: Qualitative results on PerSeg. Ground truth in green, prediction with the highest confidence in red (class names shown refer to personal class labels in PerSeg). YOLO refers to the baseline. *: refers to MOCHA (AuXFT).
@@
Table 3: Ablation studies on MOCHA. First block ablates on PCA dimension $d_{t}$. The second block changes different MOCHA components. The third block compares different pretraining: OpenImages (OI), COCO, and AuXFT.
Feature Similarity.
4.3 Main Results
We begin our evaluation in the oracle setup, where models are evaluated using ground-truth bounding boxes during the few-shot personalization stage. This setting removes the detection head and focuses purely on the classification capabilities of each approach, allowing us to compare different teacher models and types of supervision embeddings (linguistic, visual, and multimodal). Tab. 1 is structured in two blocks. The top block reports results from using teacher embeddings directly in the personalization stage, while the bottom block shows YOLO models distilled using those same supervision signals. In the first block, visual teachers show strong performance. Notably, CLIP outperforms DINO (used in AuXFT). On the other hand, LLaVa’s features ($h_{i}$) perform worse, but combining them with CLIP features ($z_{V,i}$) — i.e., using $u_{i}$ for supervision — yields notable gains at no added cost (as CLIP is already part of LLaVa), demonstrating the complementarity of visual and textual cues. Normalizing the visual features also leads to a slight improvement over their unnormalized version ($\gamma=1$). In the second block, we evaluate distillation using either visual supervision (CLIP, $z_{V,i}$), textual supervision (LLaVa, $h_{i}$), or their combination (LLaVa, $u_{i}$). All models are distilled through MOCHA supervision, showing consistent gains of our approach over the non-distilled baseline (YOLO only). Still, using a single modality alone proves less effective than leveraging full multimodal supervision.
Comparisons.
We then evaluate the complete approach, including both backbone and detection head, in the standard few-shot personalization setting (without ground-truth bounding boxes). Tab. 2 shows a comparison between MOCHA and recent approaches for personalized detection. Among distillation-based methods, KL divergence and OFA yield only marginal improvements over the YOLO baseline, while combining KL with MSE provides slightly better performance. ViLD struggles in highly personalized scenarios, despite its open-vocabulary capabilities. AuXFT achieves the strongest average performance among prior methods, but still underperforms in more challenging domains (e.g., PerSeg and POD), where fine-grained semantic alignment is critical. In contrast, MOCHA surpasses all other approaches showing consistently strong results across all datasets and settings, obtaining a +10.1 average score improvement over the YOLO baseline and +4.9 on the best competitor, AuXFT. When distilled from COCO-pretrained weights, MOCHA already exceeds the performance of state-of-the-art methods. Starting from AuXFT-pretrained weights yields further gains, confirming the compatibility and effectiveness of the initialization. Overall, the results validate the effectiveness of MOCHA in transferring high-level semantics from frozen vision-language models into compact detectors, enabling robust personalization under few-shot constraints.
Qualitative Results.
Fig. 7 presents a qualitative comparison on the PerSeg dataset, featuring three sample images: a can, a cat, and a dog. Green bounding boxes indicate ground truth annotations, while red boxes show the highest-confidence predictions. MOCHA (AuXFT) consistently demonstrates superior performance across all samples, with correct class predictions and precise bounding boxes. In contrast, competing methods exhibit occasional misclassifications and less accurate detections. The results emphasize the effectiveness of MOCHA’s feature translation and relational alignment strategies in enhancing spatial precision and adaptability to novel classes.
4.4 Ablation Studies
Tab. 3 reports an extensive ablation study, isolating the contributions of key design and training components of MOCHA across the PerSeg and iCubWorld datasets. We first assess the impact of the PCA dimensionality $d_{t}$ used to compress teacher embeddings. As previously highlighted in Fig. 4, compressing the teacher’s embedding improves performance compared to using full-resolution features. A PCA dimension of $d_{t}=512$ strikes a good balance between compactness and expressiveness. More aggressive compression (e.g., $256$ or $384$) leads to reduced performance, indicating a loss of relevant semantic content. Next, we examine the role of core MOCHA components. Removing the embedding alignment loss ($\mathcal{L}_{\mathrm{emb}}$) degrades accuracy on both datasets, underscoring its importance for preserving global teacher structure. Disabling attention mechanisms in the translator ($\mathrm{t}_{S}$ no attn.) also leads to a consistent drop, suggesting that attention layers play a key role in adapting multimodal features. Likewise, removing the translator’s final layer ($\mathrm{t}_{S}$ no ffn) reduces performance. Enabling data augmentation (standard YOLO augmentations only varying color, $\mathcal{D}_{c}$ augm.) lowers robustness, while manipulating learning rates — doubling them for encoder or decoder separately — slightly harms performance. Finally, we compare different pretraining strategies. When pretraining is performed on OpenImages only, performance remains below that of MOCHA configurations. COCO-pretrained weights yield a substantial boost, but AuXFT-pretrained weights lead to the best results overall, thanks to their strong initial performance and architectural compatibility with our MOCHA.
@@ 5 Conclusion
In this work, we introduced MOCHA, a novel approach for multimodal, cross-architecture distillation tailored to few-shot personalized object detection. MOCHA aligns compact student features to region-level embeddings derived from a large-scale frozen vision-language model, removing redundancy via PCA. Our approach leverages both intra-class and inter-class alignment, through a distillation loss and an embedding loss, respectively, achieving significant gains over the state of the art. MOCHA requires minimal overhead and operates without the teacher model at test time, making it compatible with lightweight detectors and suitable for resource-constrained applications such as mobile robotics.
Appendix
In this appendix, we provide additional experiments and analyses to validate our methodology further. We start by providing a statistical validation of MOCHA using the Wilcoxon signed-rank test to assess the significance of our findings. We continue by presenting additional results, including PCA curve fitting analyses and detailed performance breakdowns across individual datasets. We then outline the implementation details of our experiments and provide a pseudocode description of the MOCHA methodology. Finally, we conclude by presenting some qualitative results and discussing MOCHA’s limitations.
Appendix A.1 Wilcoxon Experiment
To verify the statistical significance of the results presented in the main document, in Tab. A.4 we report the Wilcoxon signed rank metric computed on the mAP achieved by MOCHA and its competitors in the PerSeg and POD datasets. The analysis is performed on the same episodic experiments in Tab. 1 (in the main document), where average scores are reported. The analysis strongly confirms the statistical superiority of MOCHA, achieving, in most cases, an extremely low $p$-value $10^{-17}$.
@@ Appendix A.1 Wilcoxon Experiment
Table A.4: $p$-value of Wilcoxon signed rank metric on PerSeg/POD 1-shot between MOCHA (AuXFT) and competitors.
Appendix A.2 Dimensionality Reduction

### v4

@@ 4.2 Analyses
Fig. 4 evaluates the role of teacher supervision dimensionality ($d_{t}$) by comparing features: (i) of the personal datasets $\mathcal{D}_{f}$ (as an upper bound), (ii) of the distillation dataset $\mathcal{D}_{c}$ unnormalized and (iii) PCA normalized (additional details and validation in Appendix). Overall, performance increases steadily up to $d_{t}=512$, which strikes a good balance between compactness and expressivity. After that, it plateaus or slightly degrades up to the maximum dimension $d=4608$. This indicates that most of the relevant semantics are preserved in a compact subspace, and that further dimensions may introduce redundancy or noise. Normalized features $u_{i}^{\prime}$ consistently outperform their unnormalized counterparts $\hat{u}_{i}^{\prime}$ (especially in the 1-shot setting), obtaining better performance even if compared to personal datasets’ features.
YOLOv8n
@@
Figure 7: Qualitative results on PerSeg. Ground truth in green, prediction with the highest confidence in red (class names shown refer to personal class labels in PerSeg). YOLOv8n refers to baseline. *: refers to MOCHA (AuXFT).
Feature Similarity.
4.3 Main Results
We begin our evaluation in the oracle setup, where models are assessed using ground truth bounding box coordinates (one per evaluation) during the few-shot personalization stage. This setting removes the detection head and focuses solely on the classification capability of each approach, allowing us to compare different teacher models and types of supervision embeddings (linguistic, visual, and multimodal). Tab. 1 is organized into two blocks. The top block reports results obtained by using the teacher embeddings directly during personalization, while the bottom block shows the performance of YOLO students distilled with the same supervision signals. In the first block, visual teachers show strong performance. Notably, CLIP outperforms DINO (used in AuXFT). On the other hand, LLaVa’s features ($h_{i}$) perform worse, but combining them with CLIP features ($z_{V,i}$) — i.e., using $u_{i}$ for supervision — yields notable gains at no added cost (as CLIP is already part of LLaVa), demonstrating the complementarity of visual and textual cues. Normalizing the visual features also leads to a slight improvement over their unnormalized version ($\gamma=1$). In the second block, we evaluate distillation using either visual supervision (CLIP, $z_{V,i}$), textual supervision (LLaVa, $h_{i}$), or their combination (LLaVa, $u_{i}$). All models are distilled through MOCHA supervision, showing consistent gains of our approach over the non-distilled baseline (YOLO only). Still, using a single modality alone proves less effective than leveraging full multimodal supervision. Note that in Tab. 1 we evaluate using the ground truth boxes, so the retrieval accuracy corresponds to standard detection accuracy, with each image containing a single annotated instance. In Tab. 2, instead, the retrieval protocol [3] allows multiple candidate boxes per image; we denote an object as correct if any predicted box matches the target. As this metric accepts multiple candidates, the scores in Tab. 2 are naturally slightly higher than those in Tab. 1.
Comparisons.
We evaluate the complete approach, including both backbone and detection head, in the standard few-shot personalization setting (predicting bounding box coordinates). Tab. 2 shows a comparison between MOCHA and recent approaches for personalized detection. Among distillation-based methods, KL divergence and OFA yield only marginal improvements over the YOLOv8n baseline, while combining KL with MSE provides slightly better performance. ViLD struggles in highly personalized scenarios, despite its open-vocabulary capabilities. AuXFT achieves the strongest average performance among prior methods, but still underperforms in more challenging domains (e.g., PerSeg and POD), where fine-grained semantic alignment is critical. In contrast, MOCHA surpasses all other approaches showing consistently strong results across all datasets and settings, obtaining a +10.1 average score improvement over the YOLOv8n baseline and +4.9 on the best competitor, AuXFT. When distilled from COCO-pretrained weights, MOCHA already exceeds the performance of state-of-the-art methods. Starting from AuXFT-pretrained weights yields further gains, confirming the compatibility and effectiveness of the initialization. MOCHA introduces only minimal computational overhead: on YOLOv8n on PerSeg, MOCHA only adds $3$ ms/image, corresponding to a $\sim$10% relative increase (further details in Appendix).
Qualitative Results.
Fig. 7 presents a qualitative comparison on the PerSeg dataset, featuring three sample images: a can, a cat, and a dog. Green bounding boxes indicate ground truth annotations, while red boxes show the highest-confidence predictions. MOCHA (AuXFT) consistently demonstrates superior performance across all samples, with correct class predictions and precise bounding boxes. In contrast, competing methods exhibit occasional misclassifications and less accurate detections. These qualitative examples reflect the trends observed in the quantitative results, with MOCHA offering more reliable predictions.
4.4 Ablation Studies
Generalization to Other Student Models.
Tab. 7 reports an ablation on the student model. Results show that MOCHA consistently improves few-shot detection across all student architectures, demonstrating strong generalization beyond the YOLOv8n model used in the main experiments. The largest gains are obtained with YOLOv8n, where the combination of MOCHA’s region-level supervision and YOLO’s multi-scale feature hierarchy yields substantial improvements over both the baseline and AuXFT. Upgrading the student to YOLOv11n preserves this trend, albeit with smaller margins. This suggests that newer backbones, while stronger on their own, leave less room for improvement but still benefit from MOCHA supervision. On the other hand, the transformer-based RT-DETR-l shows the smallest but still consistent gains. This is because RT-DETR-l is much larger ($\approx 10$ times) and architecturally very different from YOLO detectors, making the translation module less effective and limiting its suitability for on-device use. Despite this, MOCHA still provides consistent gains across datasets, confirming its robustness and ability to offer complementary supervision across detector families.
Supervision Student PerSeg iCubWorld
1 shot 1 shot 5 shot
YOLOv8n 41.2 $\pm 2.6$ 51.2 $\pm 3.0$ 68.4 $\pm 2.4$
AuXFT YOLOv8n 48.8 $\pm 3.4$ 55.0 $\pm 3.0$ 74.5 $\pm 2.5$
MOCHA YOLOv8n 59.1 $\pm 3.3$ 61.4 $\pm 3.1$ 77.0 $\pm 2.3$
YOLOv11n 49.3 $\pm 3.8$ 52.7 $\pm 4.1$ 72.8 $\pm 2.7$
AuXFT YOLOv11n 52.8 $\pm 3.8$ 53.9 $\pm 4.1$ 73.3 $\pm 2.4$
MOCHA YOLOv11n 56.2 $\pm 3.2$ 57.5 $\pm 2.5$ 74.1 $\pm 2.4$
RT-DETR-l 43.5 $\pm 3.5$ 45.0 $\pm 3.8$ 63.5 $\pm 2.5$
AuXFT RT-DETR-l 45.2 $\pm 3.4$ 46.1 $\pm 3.7$ 64.0 $\pm 2.4$
MOCHA RT-DETR-l 47.0 $\pm 3.1$ 47.8 $\pm 3.2$ 64.4 $\pm 2.3$
Table 3: Ablation studies on the student model. Comparison between AuXFT and MOCHA (AuXFT) supervisions using different student models: YOLOv8n, YOLOv11n, RT-DETR-l.
Components Design.
Tab. 4 reports an extensive ablation study, isolating the contributions of key design and training components of MOCHA across the PerSeg and iCubWorld datasets. We first assess the impact of the PCA dimensionality $d_{t}$ used to compress teacher embeddings. As previously highlighted in Fig. 4, compressing the teacher’s embedding improves performance compared to using full-resolution features. A PCA dimension of $d_{t}=512$ represents the best balance between compactness and expressiveness. For this reason, we selected it as the embedding size. More aggressive compression (e.g., $256$ or $384$) leads to reduced performance, indicating a loss of relevant semantic content, while larger ones (e.g., $1024$) increase computational cost without significant gains. Next, we examine the role of core MOCHA components. Removing the embedding alignment loss ($\mathcal{L}_{\mathrm{emb}}$) degrades accuracy on both datasets, underscoring its importance for preserving global teacher structure. Disabling attention mechanisms in the translator ($\mathrm{t}_{S}$ no attn.) also leads to a consistent drop, suggesting that attention layers play a key role in adapting multimodal features. Likewise, removing the translator’s final layer ($\mathrm{t}_{S}$ no ffn) reduces performance. Enabling data augmentation (standard YOLO augmentations only varying color, $\mathcal{D}_{c}$ augm.) lowers robustness, while manipulating learning rates — doubling them for encoder or decoder separately — slightly harms performance. Finally, we compare different pretraining strategies. When pretraining is performed on OpenImages only, performance remains below that of MOCHA configurations. COCO-pretrained weights yield a substantial boost, but AuXFT-pretrained weights lead to the best results overall, thanks to their strong initial performance and architectural compatibility with our MOCHA.
@@ 4.4 Ablation Studies
Table 4: Ablation studies on MOCHA components. First block ablates on PCA dimension $d_{t}$. The second block changes different MOCHA components. The third block compares different pretraining: OpenImages (OI), COCO, and AuXFT.
@@ 5 Conclusion
In this work, we introduced MOCHA, an effective multimodal objects-aware cross-architecture alignment technique that transfers region-level knowledge from large multimodal foundation models into compact vision detectors, enabling robust few-shot personalized object detection. Across four personal benchmarks (PerSeg, POD, CORe50, and iCubWorld), MOCHA consistently surpasses prior approaches, achieving an average improvement of +10.1 over the YOLOv8n baseline and outperforming the best competitor, AuXFT, by +4.9. Our method generalizes well across different student architectures, including YOLOv11n and RT-DETR-l. MOCHA transfers knowledge from large multimodal models into compact visual detector which are well suited for edge deployment in resource-constrained scenarios, such as mobile and robotic platforms.
Appendix A Appendix
In this appendix, we provide additional experiments and analyses to further validate our methodology. We begin with a statistical validation of MOCHA using the Wilcoxon signed-rank test to assess the significance of our findings. We then present a comprehensive performance evaluation, including additional experiments that vary either the student model or the FSL architecture. Next, we outline the implementation details of our experiments and provide pseudocode describing the MOCHA pipeline. Finally, we detail the PCA curve fitting procedure and report per-dataset performance breakdowns. We present qualitative results and discuss the current limitations of MOCHA.
A.1 Wilcoxon Experiment
To verify the statistical significance of the results presented in the main document, in Tab. 5 we report the Wilcoxon signed rank metric computed on the mAP achieved by MOCHA and its competitors in the PerSeg and POD datasets. The analysis is performed on the same episodic experiments in Tab. 2 (in the main document), where average scores are reported. The results strongly confirm the statistical superiority of MOCHA, yielding extremely small $p$-values (often below $10^{-17}$) across the majority of comparisons.
@@ A.1 Wilcoxon Experiment
Table 5: $p$-value of Wilcoxon signed-rank metric on PerSeg/POD 1-shot between MOCHA (AuXFT) and competitors.
A.2 Perfomance Evaluation
Tab. 6 summarizes the key computational properties of each supervision strategy, reporting the number of parameters, training time (in hours), inference throughput (ms/im), maximum VRAM usage, and model size (in MB). The table shows how MOCHA requires much less training time than its closest competitor, almost matching the time needed to train the detector alone. This is, in part, due to the efficient implementation of the MOCHA distillation system. Since we employ deterministic augmentation during training, it is possible to cache the mixed LLAVA-CLIP embedding vectors after the first computation. As a consequence, the overall computational complexity is reduced by several orders of magnitude, yielding a dramatic speedup over a full training. At the same time, MOCHA maintains a very lightweight inference footprint, introducing only a minimal overhead compared to AuXFT while providing significantly stronger downstream performance.
Supervision Params Pretraining FSL Inference
Dist. Complexity Time [hours] Time [ms/im] VRAM [MB] Size [MB]
3.23.2M $O(\bar{N})$ 1010h 50.5 $\pm 5.3$ $221.5$ $12.2$
AuXFT 3.93.9M $O(\bar{N}\times\bar{H}\times\bar{W})$ 5050h 32.6 $\pm 5.1$ $288.7$ $18.4$
MOCHA (AuXFT) 4.54.5M $O(\bar{N})$ 1111h 35.2 $\pm 4.6$ $352.5$ $22.8$
Table 6: Performance evaluation. Comparison of computational cost, memory usage, and inference efficiency across supervision strategies.
Supervision Student Params PerSeg POD CoRE50 iCubWorld Avg
1 shot 1 shot 5 shot 1 shot 5 shot 1 shot 5 shot
YOLOv8n 3.23.2M 41.2 $\pm 2.6$ 23.6 $\pm 2.4$ 30.4 $\pm 0.0$ 57.8 $\pm 6.3$ 67.3 $\pm 7.7$ 51.2 $\pm 3.0$ 68.4 $\pm 2.4$ 48.6
AuXFT YOLOv8n 3.93.9M 48.8 $\pm 3.4$ 31.5 $\pm 2.7$ 38.8 $\pm 0.0$ 58.8 $\pm 6.4$ 69.3 $\pm 6.0$ 55.0 $\pm 3.0$ 74.5 $\pm 2.5$ 53.8
MOCHA YOLOv8n 4.54.5M 59.1 $\pm 3.3$ 36.3 $\pm 3.3$ 45.9 $\pm 0.0$ 60.9 $\pm 4.4$ 70.6 $\pm 4.5$ 61.4 $\pm 3.1$ 77.0 $\pm 2.3$ 58.7
YOLOv11n 2.62.6M 49.3 $\pm 3.8$ 23.7 $\pm 3.0$ 30.5 $\pm 0.0$ 32.2 $\pm 5.8$ 55.3 $\pm 7.2$ 52.7 $\pm 4.1$ 72.8 $\pm 2.7$ 45.2
AuXFT YOLOv11n 3.33.3M 52.8 $\pm 3.8$ 27.3 $\pm 2.8$ 37.9 $\pm 0.0$ 33.0 $\pm 6.0$ 57.0 $\pm 6.5$ 53.9 $\pm 4.1$ 73.3 $\pm 2.4$ 47.9
MOCHA YOLOv11n 3.93.9M 56.2 $\pm 3.2$ 29.0 $\pm 2.2$ 38.2 $\pm 0.0$ 58.6 $\pm 5.9$ 66.8 $\pm 2.0$ 57.5 $\pm 2.5$ 74.1 $\pm 2.4$ 54.3
YOLOv11s 9.49.4M 50.6 $\pm 3.4$ 27.0 $\pm 2.8$ 40.4 $\pm 0.0$ 42.6 $\pm 6.2$ 63.2 $\pm 8.3$ 52.9 $\pm 3.6$ 72.4 $\pm 2.7$ 49.9
AuXFT YOLOv11s 10.110.1M 54.6 $\pm 3.4$ 32.7 $\pm 3.5$ 42.2 $\pm 0.0$ 45.2 $\pm 6.5$ 67.8 $\pm 7.0$ 54.5 $\pm 3.5$ 75.3 $\pm 2.3$ 53.2
MOCHA YOLOv11s 10.710.7M 57.0 $\pm 3.4$ 37.7 $\pm 3.1$ 46.3 $\pm 0.0$ 56.6 $\pm 6.7$ 68.6 $\pm 7.2$ 58.7 $\pm 3.0$ 76.7 $\pm 2.5$ 57.4
YOLOv11l 25.325.3M 47.6 $\pm 3.6$ 26.3 $\pm 3.0$ 36.9 $\pm 0.0$ 38.9 $\pm 7.6$ 60.5 $\pm 8.6$ 49.9 $\pm 3.7$ 65.6 $\pm 2.6$ 46.5
AuXFT YOLOv11l 26.026.0M 49.4 $\pm 3.7$ 32.4 $\pm 3.1$ 46.0 $\pm 0.0$ 40.1 $\pm 7.9$ 63.5 $\pm 8.4$ 53.1 $\pm 4.1$ 72.2 $\pm 2.4$ 51.0
MOCHA YOLOv11l 26.626.6M 56.3 $\pm 3.9$ 38.8 $\pm 2.7$ 47.6 $\pm 0.0$ 53.8 $\pm 6.3$ 61.0 $\pm 7.9$ 56.8 $\pm 3.4$ 75.0 $\pm 2.3$ 55.6
Table 7: Ablation studies on the student model. Comparison between AuXFT and MOCHA (AuXFT) supervisions using different student models: YOLOv8n, YOLOv11n, YOLOv11s, YOLOv11l.
A.3 Additional Ablation Studies
In this section we provide further ablation experiments to support our findings.
A.3.1 Student Model
In Tab. 7 we report the performance of MOCHA when applied to different YOLO architectures: YOLOv8n [17] (architecture used in the main experiments), YOLOv11n/s/l [16] (more recent and higher-performing versions). MOCHA consistently outperforms both the baseline and the closest competitor, across all settings and datasets. Interestingly, the magnitude of the improvement tends to decrease with larger models: while lightweight models such as YOLOv8n and YOLOv11n benefit from improvements exceeding 4% mIoU, the gains become more moderate for the larger YOLOv11s and YOLOv11l architectures. We hypothesize that this behaviour results from the higher representational capacity of large networks, which can reduce the effectiveness of MOCHA’s efficiency-oriented distillation strategy.
A.3.2 FSL Architecture
To further assess the robustness of MOCHA’s embeddings, Tab. 8 compares the performance of standard YOLOv8n baseline, AuXFT, and MOCHA (AuXFT) varying the architecture used for Few-Shot Learning classification. We used the same FSL architectures as in [3] plus a learned linear probing classifier. MOCHA consistently provides the best results, across all settings and datasets, demonstrating that the distilled representations remain highly effective regardless of the downstream FSL classifier. This confirms MOCHA’s ability to generalize reliably across different learning paradigms, from metric-based approaches such as ProtoNet [38] and SimpleShot [43] to parametric classifiers such as Linear Probing.
Supervision FSL Architecture PerSeg POD
1 shot 1 shot
ProtoNet (AuXFT) 41.2 $\pm 2.6$ 23.6 $\pm 2.4$
AuXFT ProtoNet (AuXFT) 48.8 $\pm 3.4$ 31.5 $\pm 2.7$
MOCHA ProtoNet (AuXFT) 59.1 $\pm 3.3$ 36.3 $\pm 3.3$
ProtoNet ($\ell_{2}$) 43.4 $\pm 3.4$ 16.5 $\pm 2.3$
AuXFT ProtoNet ($\ell_{2}$) 54.1 $\pm 3.4$ 31.0 $\pm 2.9$
MOCHA ProtoNet ($\ell_{2}$) 62.7 $\pm 3.0$ 32.7 $\pm 2.9$
SimpleShot 42.8 $\pm 3.1$ 16.4 $\pm 2.4$
AuXFT SimpleShot 53.7 $\pm 3.2$ 30.5 $\pm 2.8$
MOCHA SimpleShot 61.6 $\pm 3.4$ 33.3 $\pm 2.8$
Linear Probing  2.9 $\pm 1.2$ 1.5 $\pm 1.0$
AuXFT Linear Probing 22.8 $\pm 2.9$ 9.6 $\pm 2.0$
MOCHA Linear Probing 42.8 $\pm 3.0$ 19.5 $\pm 2.4$
Table 8: Ablation studies on the FSL architecture. Comparison between AuXFT and MOCHA (AuXFT) supervisions using different Few-Shot Learners: ProtoNet (AuxFT), ProtoNet, SimpleShot, Linear Probing (single linear with bias; Adam, $lr=0.1$).
A.3.3 Dimensionality Reduction